# When People are Floods: Analyzing Dehumanizing Metaphors in Immigration Discourse with Large Language Models

### v1

@@ 7.1 Ideology’s Role in Metaphor
Dependent variables are metaphor scores for each concept. Fixed effects include a binary ideology score and a continuous ideology strength score (group-mean centered and z-score normalized), and the interaction between these two variables. These variables are derived from the original ideology estimates from Mendelsohn et al. (2021). Breaking this score to separately account for direction (ideology) and magnitude (ideology strength) allows us to draw conclusions about the far-left, far-right, and moderates.
We control for message, author, and time variables (e.g., tweet length, follower count, year and month) as fixed effects. For robustness, we additionally specify models that control for frames already included in the Immigration Tweets Dataset. [^6] We control for ten issue-generic policy frames
@@
See regression tables (A8-A9) for all included variables. We assess statistical significance at the $p=0.05$ level after applying Holm-Bonferroni corrections to account for multiple comparisons.
Figure 4: Marginal effect of conservative ideology on metaphor scores, estimated from regression models.
Figure 5: Marginal effect of ideology strength on metaphor scores, estimated from regression models.
@@ Results
Addressing RQ2, the relationship between ideology strength and metaphor differs for liberals and conservatives (Fig. 5). Among conservatives, extremism is associated with more metaphor across all concepts. For liberals, however, the relationship between strength and metaphor depends on the concept. Extreme liberal ideology is associated with lower use of water and commodity but higher use of creature-related metaphors (parasite, vermin, animal). In sum, both extremes are associated with greater use of creature-related metaphor, but only stronger conservative ideology is associated with greater use of object-related metaphor. These findings hold when controlling for topical frames (Appendix Table A9 and Fig. A7).
@@ 7.2 Metaphor’s Role in Engagement
We fit linear regression models, where dependent variables are favorite and retweet counts (log-scaled). Independent variables include all concepts’ metaphor scores, ideology, ideology strength, and interactions between scores and ideology. We control for the same variables as in §7.1 (including verified status and follower counts, which strongly predict engagement). We also fit models controlling for topic-like frames.
Figure 6: Average marginal effect of metaphor scores on retweets (log-scaled) for each concept.
Figure 7: Group-average marginal effects of metaphor on (log-scaled) retweets separated by ideology.
@@ Results
Aligned with prior evidence that source domains moderate metaphors’ effects (Bosman, 1987), associations between metaphor scores and user engagement vary by concept (Fig. 6). Creature metaphors (vermin,parasite, and animal) are associated with more retweets, with the largest effects for liberals (Fig. 7). Only parasite is significantly associated with more favorites, and metaphors from some domains (e.g. commodity) are actually associated with fewer favorites (Appendix Fig. A8). For both favorites and retweets, the direction of the effect diverges for only one concept: water, which is associated with higher engagement for conservatives, but lower for liberals. See Appendix Figs. A9-A10 for results when controlling for frames, and Tables A10-A11 for full regression coefficients.
@@ 7.3 How are liberals using metaphors?
Our quantitative analysis paints a complex picture for liberals’ use of dehumanizing metaphors of immigrants.While conservative ideology is more associated with using these metaphors (§7.1), liberals use them to a substantial extent (Appendix Fig. A5). Furthermore, more extreme liberal ideology is associated with greater use of creature-related metaphors, and the positive relationship between creature-related metaphors and higher retweet counts is driven by liberals. We thus conduct a qualitative analysis of 25 liberal tweets with the highest scores for each concept. Examples discussed here are shown in Appendix Table A7.
We identify four themes: (i.) Straightforwardly embracing metaphors. Liberals describe migrants as a source of \$\$\$ (commodity) and a wave (water). (ii.) Sympathetic framing, particularly to highlight humanitarian concerns. For example, a liberal overtly cues the animal metaphor while lamenting that “they hunt them like animals, they cage them like animals”. Other sympathetic instances refer to feeding and sheltering immigrants; while these verbs can be used to talk about humans, they deny agency to the recipient and are thus more associated with animals (Tipler and Ruscher, 2014). (iii.) Quoting or paraphrasing outpartisans to criticize their use of inflammatory metaphors. (e.g., paraphrasing conservative politicians who called immigrants “rats” and an “infestation”.) (iv.) Redirecting dehumanizing metaphors from immigrants to political opponents. For example talking about “right-wing forces” (war metaphor) or referring to Melania Trump (Donald Trump’s wife and an immigrant) as a “tick” and “blood sucker”.
@@ 8 Discussion
Our work offers many avenues for future research. Future work could adapt and evaluate our methodology for other issues, languages, and cultures. While we curate source domains from social science literature, such resources may not be available for lesser-studied contexts. Future work could thus explore developing automated methods for metaphor discovery, possibly with the aid of external knowledge graphs and lexical resources (Mao et al., 2022, 2023). Future analysis-oriented work could use NLP-based measurements of metaphorical language in large-scale experiments to precisely quantify the effects of metaphor on emotions, policy preferences, and social attitudes.
@@ 10 Ethical Implications
Our primary ethical concerns relate to our reporting of dehumanizing metaphors. Even though we clearly do not endorse these dehumanizing metaphors, merely exposing them to annotators and readers risks reinforcing harmful conceptual associations. Merely reporting others’ use of slurs can still harm members of targeted communities (Croom, 2011). It remains an open question if reporting dehumanizing metaphor (even to vehemently disagree with their premise) has similar effects as straightforward usage.
We recruited hundreds of annotators to help us create our evaluation dataset. The study was deemed exempt by the Anonymized University Institutional Review Board and annotators were fairly compensated at an average rate of \$16/hour. Nevertheless, creating this dataset involved exposing annotators to offensive and hateful social media posts. We attempt to mitigate these harms by flagging the task as sensitive on Prolific, warning participants of its potentially harmful nature, and limiting each task to just 20 tweets, of which only a few are typically overtly hateful.
@@ Appendix A Appendix
This appendix contains additional details about our annotation process (§A.1), methodology (§A.2), model evaluation (§A.3), and analysis (§A.4).
@@ A.3 Evaluation
A.4 Analysis
A.4.1 Descriptive Analysis of Scores
This section includes descriptive analyses of metaphor scores (§A.4.1), results for all regression-based analyses (§A.4.2, A.4.3), and example tweets with high metaphor scores (Table A7).
@@ A.4.1 Descriptive Analysis of Scores
Figure A7: Average marginal effect of conservative ideology (left) and ideological strength (right) on metaphor. Effects are estimated from linear regression models that control for issue-generic policy frames as fixed effects.
A.4.2 Role of Ideology in Metaphor
Figures A7 shows average marginal effects from regression models that include issue-generic policy frames as fixed effects (Mendelsohn et al., 2021). This regression also facilitates analysis of the relationships between issue-generic policy frames, metaphor, and ideology (Figure A6). Aligning with expectations, some issue-generic frames are strongly associated with particular metaphorical concepts (e.g., economic for commodity and security for war), and metaphors are more readily used with some frames compared to others (e.g., security is more metaphorical than cultural identity). Tables A8 and A9 show full regression coefficients from models that exclude and include issue-generic frames, respectively.
A.4.3 Role of Metaphor in Engagement
Figures A8 and A9 shows the average marginal effect of metaphor on favoriting behavior, excluding and including issue-generic frames as fixed effects respectively. Figure A10 illustrates average marginal effects for retweets, controlling for frames. All regression coefficients for both sets of models can be found in Tables A10-A11.
Figure A8: Average marginal effect of metaphor on favorites, from regression without issue-generic frames.
Figure A9: Average marginal effect of metaphor on favorites, from regression including issue-generic frames.
Figure A10: Average marginal effect of metaphor on retweets, from regression including issue-generic frames.

### v2

@@ 7.1 Ideology’s Role in Metaphor
Dependent variables are metaphor scores for each concept. Fixed effects include a binary ideology variable (liberal or conservative) and a continuous ideology strength score (group-mean centered and z-score normalized), and the interaction between these two variables. These variables capture distinct components of political ideology and enable us to draw conclusions about the far-left, far-right, and moderates.
Figure 4: Effect of conservative ideology (relative to liberal ideology) on metaphor score, shown as the average percent change. Percent changes are calculated from marginal effects estimated from regression models.
We control for message, author, and time variables (e.g., tweet length, follower count, year and month) as fixed effects. For robustness, we also specify models that control for frames included in the Immigration Tweets Dataset. [^6] We control for ten issue-generic policy frames
@@
See regression tables (A8-A9) for all included variables. We assess significance at the $p=0.05$ level after applying Holm-Bonferroni corrections to account for multiple comparisons. Due to the presence of interaction terms, we enhance interpretability by calculating average marginal effects and visualizing effects as the average percent change in metaphor score.
Figure 5: Effect of a one standard deviation increase in ideological strength on metaphor scores. Percent changes are calculated from group-average marginal effects estimated from regression models. Group-average marginal effects reveal that ideological strength has distinct effects on metaphor use among liberals and conservatives, which further vary across concepts.
@@ Results
Addressing RQ2, the relationship between ideology strength and metaphor differs for liberals and conservatives (Fig. 5). Among conservatives, extremism is associated with more metaphor across all concepts. For liberals, however, the relationship between strength and metaphor depends on the concept. Extreme liberal ideology is associated with lower use of water and commodity but higher use of creature-related metaphors (parasite, vermin, animal). In sum, both extremes are associated with greater use of creature-related metaphor, but only stronger conservative ideology is associated with greater use of object-related metaphor. These findings hold when controlling for topical frames (Table A9 and Figs. A7-A8).
@@ 7.2 Metaphor’s Role in Engagement
We fit linear regression models, where dependent variables are favorite and retweet counts, log-transformed as $ln(x+1)$. Independent variables include all concepts’ metaphor scores, ideology, ideology strength, and interactions between scores and ideology. We control for the same variables as in §7.1 (including verified status and follower counts, which strongly predict engagement). We again additionally fit models controlling for issue-generic frames.
Interpreting the resulting regression coefficients directly is challenging. As in §7.1, we calculate average marginal effects, and further back-transform estimates from the $ln(x+1)$ scale to get percent changes in retweet counts. Although the metaphor score is standardized, it is an abstract measure for which one standard deviation difference is not easily interpretable on its own. We thus focus on a ±2 standard deviation range to capture contrasts between tweets with very low and high metaphorical content, providing a more intuitive sense of how metaphorical language relates to user engagement.
Figure 6: Average marginal effect of metaphor scores on retweets. Effects are shown as percent changes in predicted retweets between non-metaphorical and highly-metaphorical tweets (±2 standard deviations).
Figure 7: Group-average marginal effects of metaphor on retweets by ideology. Percent changes in predicted retweets between non-metaphorical and highly-metaphorical tweets (±2 standard deviations) are shown.
@@ Results
Aligned with prior evidence that source domains moderate metaphors’ effects (Bosman, 1987), associations between metaphor scores and user engagement vary by source domain (Fig. 6). Creature metaphors (vermin, parasite, and animal) are associated with more retweets, with the largest effects among liberals (Fig. 7). Only parasite is significantly associated with more favorites, and metaphors from some domains (e.g. commodity) are actually associated with fewer favorites (Appendix Figs. A9-A10). For both favorites and retweets, the direction of the effect diverges for only one concept: water, which is associated with higher engagement for conservatives, but lower for liberals. See Appendix Figs. A11-A12 for results when controlling for frames, Tables A10-A11 for full regression coefficients, and Tables A12-A13 for marginal effects estimates.
@@ 7.3 How are liberals using metaphors?
Our quantitative analysis paints a complex picture for liberals’ use of dehumanizing metaphors of immigrants. While conservative ideology is associated with higher metaphor usage across source domains (§7.1), liberals use them to a substantial extent (Appendix Fig. A5). Furthermore, among liberals, more extreme ideology is associated with higher creature-related metaphor scores, and the positive relationship between creature-related metaphors and higher retweet counts is driven by liberals. We thus conduct a qualitative analysis of 25 liberal tweets with the highest scores for each concept. Examples discussed here are shown in Table A7. We identify four themes:
  1. 1.
    Straightforwardly embracing metaphors. Liberals describe migrants as a source of \$\$\$ (commodity) and a wave (water), even if they take pro-immigrant stances.
  2. 2.
    Sympathetic framing, particularly to highlight humanitarian concerns. For example, a liberal overtly cues the animal metaphor while lamenting that “they hunt them like animals, they cage them like animals”. Other sympathetic instances refer to feeding and sheltering immigrants; while these verbs can be used to talk about humans, they deny agency to the recipient and are thus more associated with animals (Tipler and Ruscher, 2014).
  3. 3.
    Quoting or paraphrasing outpartisans to highlight their use of inflammatory metaphors. For example, liberal tweets with high vermin metaphor scores quote conservative politicians who refer to immigrants as “rats” or an “infestation”, and such tweets are often critical of blatantly dehumanizing metaphors.
  4. 4.
    Redirecting dehumanizing metaphors from immigrants to outpartisans. For example, liberal tweets redirect the war metaphor from immigrants to “right-wing forces” or invoke the parasite metaphor when referring to Melania Trump (Donald Trump’s wife and an immigrant) as a “tick” and “blood sucker”.
@@ 8 Discussion
Our work offers many avenues for future research. Future work could adapt and evaluate our methodology for other issues, languages, and cultures. While we curate source domains from social science literature, such resources may not be available for lesser-studied contexts. Future research could explore developing automated methods for metaphor discovery, possibly with the aid of external knowledge graphs and lexical resources (Mao et al., 2022, 2023). Future analysis-oriented work could use NLP-based measurements of metaphorical language in large-scale experiments to precisely quantify the effects of metaphor on emotions, policy preferences, and social attitudes.
@@ 10 Ethical Implications
Our primary ethical concerns relate to our reporting of dehumanizing metaphors. Even though we clearly do not endorse these dehumanizing metaphors, merely exposing them to annotators and readers risks reinforcing harmful conceptual associations. Even the act of reporting others’ use of slurs can still harm members of targeted communities (Croom, 2011). It remains an open question if reporting dehumanizing metaphor (even to vehemently disagree with their premise) has similar effects as straightforward usage.
We recruited hundreds of annotators to help us create our evaluation dataset. The study was deemed exempt by the University of Michigan Institutional Review Board (#HUM00253235) and annotators were fairly compensated at an average rate of \$16/hour. Nevertheless, creating this dataset involved exposing annotators to offensive and hateful social media posts. We attempt to mitigate these harms by flagging the task as sensitive on Prolific, warning participants of its potentially harmful nature, and limiting each task to just 20 tweets, of which only a few are typically overtly hateful.
11 Acknowledgments
We would like to thank Dallas Card, Yulia Tsvetkov, Nicholas Valentino, audiences at MPSA and COMPTEXT 2025, and the ARR reviewers and area chair for their helpful feedback and suggestions. J.M. gratefully acknowledges support from the Google PhD Fellowship.
@@ Appendix A Appendix
This appendix contains additional details about our annotation process (§A.1), prompts, model evaluation (§A.2), and analysis (§A.3).
@@ A.2 Evaluation
A.3 Analysis
A.3.1 Descriptive Analysis of Scores
This section includes descriptive analyses of metaphor scores (§A.3.1), results for all regression-based analyses (§A.3.2, A.3.3), and example tweets with high metaphor scores (Table A7).
@@ A.3.1 Descriptive Analysis of Scores
Figure A7: Effect of conservative ideology (relative to liberal) on metaphor score, shown as the average percent change. Percent changes are calculated from marginal effects estimated from regression models that control for issue-generic policy frames as fixed effects.
A.3.2 Role of Ideology in Metaphor
Figures A7 and A8 shows average marginal effects from regression models that include issue-generic policy frames as fixed effects (Mendelsohn et al., 2021). This regression also facilitates analysis of the relationships between issue-generic policy frames, metaphor, and ideology (Figure A6). Aligning with expectations, some issue-generic frames are strongly associated with particular metaphorical concepts (e.g., economic for commodity and security for war), and metaphors are more readily used with some frames compared to others (e.g., security is more metaphorical than cultural identity). Tables A8 and A9 show full regression coefficients from models that exclude and include issue-generic frames, respectively.
Figure A8: Effect of a one standard deviation increase in ideological strength on metaphor scores. Percent changes are calculated from group-average marginal effects estimated from regression models that control for issue-generic policy frames as fixed effects.
@@ F Statistic 1,887∗∗∗ 1,369∗∗∗ 2,375∗∗∗ 1,236∗∗∗ 2,637∗∗∗ 2,365∗∗∗ 1,129∗∗∗ 1,686∗∗∗
A.3.3 Role of Metaphor in Engagement
Figures A9 and A10 shows the percent change in favorites associated with a 4 standard deviation change in metaphor score, which corresponds to the difference between a non-metaphorical and a highly-metaphorical tweet. Figure A11 illustrates average marginal effects of metaphor on retweets, controlling for issue-generic frames as fixed effects. Figure A12 separates effects on retweets by ideology. All regression coefficients for both sets of models can be found in Tables A10-A11.
Figure A9: Average marginal effect of metaphor scores on favorites for each concept, estimated from regression models. Effects are shown as percent changes in predicted favorite counts between non-metaphorical and highly-metaphorical tweets (±2 standard deviations).
Figure A10: Group-average marginal effects of metaphor on favorites separated by ideology, estimated from regression models. Effects are shown as percent changes in predicted favorite counts between non-metaphorical and highly-metaphorical tweets (±2 standard deviations).
Figure A11: Average marginal effect of metaphor scores on retweets for each concept, estimated from regression models that control for issue-generic frames. Effects are shown as percent changes in predicted retweet counts between non-metaphorical and highly-metaphorical tweets (±2 standard deviations).
Figure A12: Group-average marginal effects of metaphor on retweets separated by ideology, estimated from regression models that control for issue-generic frames. Effects are shown as percent changes in predicted retweet counts between non-metaphorical and highly-metaphorical tweets (±2 standard deviations).
@@ F Statistic 4,114∗∗∗ 4,517∗∗∗
Outcome Concept
Marginal Effect
Percent Change
Marginal Effect
Percent Change
Marginal Effect
Percent Change
retweets animal 0.005 (0.002) 0.479 (0.151) 0.001 (0.002) 0.130 (0.185) 0.010 (0.003) 1.016 (0.257)
commodity -0.004 (0.001) -0.447 (0.127) -0.004 (0.001) -0.386 (0.147) -0.005 (0.002) -0.541 (0.231)
parasite 0.009 (0.002) 0.875 (0.172) 0.004 (0.002) 0.418 (0.190) 0.016 (0.003) 1.582 (0.326)
pressure 0.003 (0.001) 0.253 (0.136) 0.001 (0.001) 0.054 (0.148) 0.006 (0.003) 0.560 (0.259)
vermin 0.010 (0.002) 0.981 (0.182) 0.008 (0.002) 0.823 (0.220) 0.012 (0.003) 1.225 (0.314)
war 0.001 (0.001) 0.099 (0.130) 0.001 (0.001) 0.144 (0.148) 0.000 (0.002) 0.029 (0.237)
water -0.002 (0.001) -0.244 (0.138) 0.004 (0.001) 0.393 (0.148) -0.012 (0.003) -1.214 (0.263)
favorites animal -0.003 (0.002) -0.305 (0.186) -0.002 (0.002) -0.167 (0.229) -0.005 (0.003) -0.517 (0.315)
commodity -0.009 (0.002) -0.917 (0.158) -0.005 (0.002) -0.525 (0.182) -0.015 (0.003) -1.517 (0.284)
parasite 0.017 (0.002) 1.751 (0.216) 0.004 (0.002) 0.381 (0.236) 0.038 (0.004) 3.892 (0.414)
pressure -0.004 (0.002) -0.382 (0.168) -0.001 (0.002) -0.144 (0.184) -0.007 (0.003) -0.746 (0.318)
vermin -0.002 (0.002) -0.151 (0.224) 0.001 (0.003) 0.060 (0.272) -0.005 (0.004) -0.473 (0.384)
war -0.007 (0.002) -0.688 (0.160) -0.002 (0.002) -0.195 (0.184) -0.015 (0.003) -1.441 (0.290)
water 0.001 (0.002) 0.059 (0.172) 0.008 (0.002) 0.780 (0.185) -0.010 (0.003) -1.037 (0.328)
Table A12: Marginal effects of one standard deviation increase in metaphor scores on user engagement, estimated from regression models. Effects for both favorites and retweets are shown, for all tweets as well as group-average marginal effects for liberals and conservatives. Marginal effects correspond to changes in the $ln(x+1)$ scale. Corresponding percent changes in user engagement counts are also shown. Standard errors are displayed in parentheses.
Outcome Concept
Marginal Effect
Percent Change
Marginal Effect
Percent Change
Marginal Effect
Percent Change
retweets animal 0.002 (0.002) 0.188 (0.153) -0.001 (0.002) -0.059 (0.186) 0.006 (0.003) 0.569 (0.258)
commodity -0.003 (0.001) -0.265 (0.132) -0.003 (0.001) -0.269 (0.150) -0.003 (0.002) -0.260 (0.235)
parasite 0.006 (0.002) 0.638 (0.175) 0.002 (0.002) 0.193 (0.191) 0.013 (0.003) 1.326 (0.327)
pressure 0.002 (0.001) 0.184 (0.138) 0.000 (0.002) 0.032 (0.150) 0.004 (0.003) 0.418 (0.261)
vermin 0.010 (0.002) 0.982 (0.184) 0.008 (0.002) 0.841 (0.222) 0.012 (0.003) 1.199 (0.315)
war 0.001 (0.001) 0.123 (0.134) 0.003 (0.002) 0.256 (0.153) -0.001 (0.002) -0.080 (0.238)
water 0.000 (0.001) 0.032 (0.140) 0.006 (0.001) 0.605 (0.150) -0.008 (0.003) -0.842 (0.265)
favorites animal -0.003 (0.002) -0.259 (0.189) -0.001 (0.002) -0.059 (0.231) -0.006 (0.003) -0.564 (0.318)
commodity -0.010 (0.002) -1.004 (0.163) -0.007 (0.002) -0.689 (0.185) -0.015 (0.003) -1.484 (0.288)
parasite 0.013 (0.002) 1.339 (0.219) 0.001 (0.002) 0.097 (0.238) 0.032 (0.004) 3.277 (0.415)
pressure -0.002 (0.002) -0.199 (0.172) 0.000 (0.002) -0.004 (0.187) -0.005 (0.003) -0.496 (0.322)
vermin 0.000 (0.002) 0.017 (0.227) 0.002 (0.003) 0.241 (0.275) -0.003 (0.004) -0.327 (0.386)
war -0.005 (0.002) -0.526 (0.166) 0.000 (0.002) 0.021 (0.190) -0.014 (0.003) -1.361 (0.293)
water 0.003 (0.002) 0.314 (0.175) 0.009 (0.002) 0.933 (0.187) -0.006 (0.003) -0.630 (0.331)
Table A13: Marginal effects of one standard deviation increase in metaphor scores on user engagement, estimated from regression models that additionally control for issue-generic frames. Effects for both favorites and retweets are shown, for all tweets as well as group-average marginal effects for liberals and conservatives. Marginal effects correspond to changes in the $ln(x+1)$ scale. Corresponding percent changes in user engagement counts are also shown. Standard errors are displayed in parentheses.